OpenStudy (anonymous):

What is the sum of the geometric series 2^0+2^1+2^2+2^3+2^4+...+2^9 ?

OpenStudy (anonymous):

there is an easy enough formula for this do you know it?

OpenStudy (anonymous):

it is even easier with powers of two than with other powers \[2^0+2^1+2^2+...+2^n=2^{n+1}-1\]

OpenStudy (anonymous):

you can divide \(\large\color{#000000 }{ \displaystyle r^{n+1}-1 }\) by \(\large\color{#000000 }{ \displaystyle r-1 }\) and you will the series

OpenStudy (anonymous):

Thank you😁

OpenStudy (anonymous):

For example \(\large\color{#000000 }{ \displaystyle 1+r+r^2+r^3+r^4=(r^{5}-1)/(r-1) }\)